A coil having inductance 2.0 H and resistance 20 Ω is connected to a battery of emf 4.0 V. Find (a) the current at the instant 0.20 s after the connection is made and (b) the magnetic field energy at this instant.

Grade:10

1 Answers

Deepak Patra
9 years ago
Sol. L = 2.0 H, R = 20 Ω, emf = 4.0 V, t = 0.20 S i base 0 = e/R = 4/20, t = L/R = 2/20 = 0.1 (a) i = i base 0 (1 – e^-0.2/0.1) = 0.17 A (b) ½ Li^2 = ½ x 2 x (0.17)^2 = 0.0289 = 0.03 J.